# FLOSS: Federated Learning with Opt-Out and Straggler Support

### v1

@@ Abstract.
doi: conference: Workshop on Practical Adoption Challenges of ML for Systems; Seoul, South Korea;
@@ 2. Notation and Problem Setup
[…]
@@
In a typical FL setup, we have a set of $n$ users $\operatorname{\mathcal{U}}$, each with their own private dataset consisting of multiple realizations of the features $X$ and outcome $Y$. A model $h_{\theta}:x\mapsto y$ is then trained in a decentralized manner to minimize the expected loss $E[L(x,y,\theta)]$, often approximated by the empirical risk $\frac{1}{n}\sum_{i=1}^{n}L(x_{i},y_{i},\theta)$, over multiple rounds. Moving forward, we suppress dependence of the loss and gradient functions on the data $x,y$ for brevity. At each step $t$, the central server samples a subset of users $\operatorname{\mathcal{U}}^{\prime}$ of size $k$, and requests gradients $G(\theta^{(t)})$ of the loss function evaluated over their private datasets. Each sampled device uploads their gradients $g_{1}(\theta^{(t)}),\dots,g_{k}(\theta^{(t)})$, or noisy and clipped versions of them to add differential privacy (Abadi et al., 2016; Birrell et al., 2024; Das et al., 2025). The central server then aggregates these gradients to obtain $\overline{g}(\theta^{(t)})$ and updates the model as $\theta^{(t+1)}\leftarrow\theta^{(t)}-\eta\cdot\overline{g}(\theta^{(t)})$, where $\eta$ is the learning rate. Note we assume equal-sized datasets for brevity of notation, but our methodology generalizes in a straightforward manner. Finally, the updated model $h_{\theta^{(t+1)}}$ is broadcast to all users $u\in{\operatorname{\mathcal{U}}}$, and the process repeats.
@@
Two key issues arise in the above process that lead to missing data: (i) Some devices, known as stragglers (Albrecht et al., 2006; Dean and Ghemawat, 2004), may fail to upload their gradients in a reasonable time frame and the server is forced to perform the aggregation step without them, and (ii) certain users may decline to share their data for training as part of a data-sharing agreement, so not all devices can be prompted for their gradients. Both of these issues can lead to systematic, rather than completely random, missingness of gradients, resulting in degraded accuracy of the final learned model if not appropriately accounted for. In the following, we formalize the kind of data-sharing agreements we support in FLOSS, which reflects mandatory user privacy agreements that are ubiquitous across modern machine learning applications (Das et al., 2025).
@@ 3. A Formal Model of Missing Data in FL
The absence of edges in an m-DAG also encode statistical relations between the variables via the well-known d-separation criterion (Pearl, 1988) defined as follows. A path in an m-DAG ${\mathcal{G}}$ is an alternating sequence of vertices and edges $V_{1}{\ \bf-\ }V_{2}{\ \bf-\ }V_{3},\dots,V_{K}$, where each “--” in the sequence is an edge $V_{k}\leftarrow V_{k+1}$ or $V_{k}\rightarrow V_{k+1}$ that exists in ${\mathcal{G}}$, and every vertex and edge in the sequence appears at most once. A vertex $V_{k}$ is said to be a collider on the path if the preceding and succeeding edges both point into it, i.e., the path contains $V_{k-1}\rightarrow V_{k}\leftarrow V_{k+1}$. Given disjoint sets of vertices $A,B$ and $C$, the sets $A$ and $B$ are said to be d-separated given $C$, denoted $A\perp\!\!\!\perp_{\text{d-sep}}B\mid C$, if and only there is no path from a vertex in $A$ to one in $B$ along which (i) every collider on the path is either in $C$ or has a descendant in $C$ and (ii) every non-collider on the path is not in $C$. Paths satisfying conditions (i) and (ii) are said to be open. This definition leads to the following global Markov property of m-DAGs—given disjoint sets $A,B,C$ we have, $A\perp\!\!\!\perp_{\text{d-sep}}B\mid C\implies A\perp\!\!\!\perp B\mid C\text{ in }p(V)$. That is, d-separation in ${\mathcal{G}}$ implies conditional independences in the probability distribution on the random variables displayed in the m-DAG. This gives us an intuitive way to reason about missingness in FL systems, as we now demonstrate.
@@
However, there is another complication arising from user opt-out that likely results in gradients that are missing not at random (MNAR). The data are MNAR if missingness is not independent of the missing variable given observed covariates alone, i.e., $R\not\perp\!\!\!\perp G(\theta)^{miss}\mid D$ in our FL setup. This occurs when user opt-out is influenced by the data $X,Y$ itself—e.g., a user may not want to share interactions with the model $h_{\theta}$ involving sensitive data $X$ or if they are dissatisfied with model predictions of their outcomes $Y$—encoded by the edges $X\rightarrow R\leftarrow Y$. This leads to open paths—e.g., $R\leftarrow Y{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\implies}G(\theta)^{miss}$—that imply $R\not\perp\!\!\!\perp G(\theta)^{miss}\mid D$.
@@ 4. Reweighted Device Selection
4:  Prompt all users $u\in\operatorname{\mathcal{U}}$ for participation, record $R$
5:  Prompt all users $u\in\operatorname{\mathcal{U}}$ for satisfaction, record $S^{miss}$
@@
7:  Define $\operatorname{\mathcal{U}}_{R}$ as users $u\in\operatorname{\mathcal{U}}$ such that $R=1$
8:  for $i$ from 111 to $max\ iterations$ do
@@
16:    Broadcast $\theta^{(t+1)}$ to all users $u\in\operatorname{\mathcal{U}}$
@@ 5. Preliminary Results & Discussion
Conclusion

### v2

@@ Abstract.
federated learning, missing data, privacy
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Practical Adoption Challenges of ML for Systems; October 13–16, 2025; Seoul, Republic of Koreabooktitle: Practical Adoption Challenges of ML for Systems (PACMI ’25), October 13–16, 2025, Seoul, Republic of Koreadoi: 10.1145/3766882.3767180isbn: 979-8-4007-2205-9/2025/10ccs: Security and privacy Privacy protectionsccs: Computing methodologies Learning settingsccs: Computing methodologies Cooperation and coordination
@@ 2. Notation and Problem Setup
$X:\text{a set of features used for generating predictions}.$
$Y:\text{the outcome of interest (real-valued or categorical)}.$
@@
$S:\text{user satisfaction with system and model performance}.$
$R:\text{binary indicator of responsiveness to server requests}$
 $\text{\hskip 11.38092pt $R=0$ for stragglers/users opting out; $R=1$ otherwise}.$
@@
In a typical FL setup, we have a set of $n$ users $\mathcal{U}$, each with their own private dataset consisting of multiple realizations of the features $X$ and outcome $Y$. A model $h_{\theta}:x\mapsto y$ is then trained in a decentralized manner to minimize the expected loss $E[L(x,y,\theta)]$, often approximated by the empirical risk $\frac{1}{n}\sum_{i=1}^{n}L(x_{i},y_{i},\theta)$, over multiple rounds. Moving forward, we suppress dependence of the loss and gradient functions on the data $x,y$ for brevity. At each step $t$, the central server samples a subset of users $\mathcal{U}^{\prime}$ of size $k$, and requests gradients $G(\theta^{(t)})$ of the loss function evaluated over their private datasets. Each sampled device uploads their gradients $g_{1}(\theta^{(t)}),\dots,g_{k}(\theta^{(t)})$, or noisy and clipped versions of them to add differential privacy (Abadi et al., 2016; Birrell et al., 2024; Das et al., 2025). The central server then aggregates these gradients to obtain $\overline{g}(\theta^{(t)})$ and updates the model as $\theta^{(t+1)}\leftarrow\theta^{(t)}-\eta\cdot\overline{g}(\theta^{(t)})$, where $\eta$ is the learning rate. Note we assume equal-sized datasets (i.e., $X$ and $Y$ are understood to contain an equal number of realizations of features and outcomes for each user) for brevity of notation, but our methodology generalizes in a straightforward manner. Finally, the updated model $h_{\theta^{(t+1)}}$ is broadcast to all users $u\in{\mathcal{U}}$, and the process repeats.
@@
Two key issues arise in the above process that lead to missing data: (i) Some devices, known as stragglers (Albrecht et al., 2006; Dean and Ghemawat, 2004), may fail to upload their gradients in a reasonable time frame and the server is forced to perform the aggregation step without them, and (ii) certain users may decline to share their data for training as part of a data-sharing agreement, so not all devices can be prompted for their gradients. Both of these issues can lead to systematic, rather than completely random, missingness of gradients.
When the data are systematically missing, performing gradient aggregation based on just the observed data leads to biased estimates relative to the aggregate that would be obtained had everyone been able and willing to share their data (possibly contrary to fact). That is, we use the term bias here in a statistical sense—a tendency of parameter and gradient estimates computed from just observed data to deviate from the values that would be obtained had there been no missing data at all. This kind of bias arising due to missing data (and the closely related problem of selection bias) is well documented in the causal inference literature, as it often leads to biased estimation of causal effects (Mohan et al., 2013; Nabi et al., 2025). On the other hand, while there has been preliminary work on missing data in FL systems when the data are missing completely at random (Valdeira et al., 2024), to our knowledge, the issues that arise in FL systems due to systematic missingness—and how to mitigate such issues—have not been studied thus far. We note also that bias due to missing data and bias in the sense of whether an algorithm is fair (or not) are, in fact, related problems; it has also been shown that mitigating the effects of missing data can also lead to improved fairness (Kallus and Zhou, 2018; Goel et al., 2021). Thus, in this work we focus our attention on mitigating bias due to missing data, with the understanding that this can, in some cases, also lead to improved performance on certain measures of algorithmic fairness.
In the following, we formalize the kind of data-sharing agreements we support in FLOSS, which reflects mandatory user privacy agreements that are ubiquitous across modern machine learning applications (Das et al., 2025).
@@ 3. A Formal Model of Missing Data in FL
The absence of edges in an m-DAG also encode statistical relations between the variables via the well-known d-separation criterion (Pearl, 1988) defined as follows. A path in an m-DAG $\mathcal{G}$ is an alternating sequence of vertices and edges $V_{1}{\ \bf-\ }V_{2}{\ \bf-\ }V_{3},\dots,V_{K}$, where each “-” in the sequence is an edge $V_{k}\leftarrow V_{k+1}$ or $V_{k}\rightarrow V_{k+1}$ that exists in $\mathcal{G}$, and every vertex and edge in the sequence appears at most once. A vertex $V_{k}$ is said to be a collider on the path if the preceding and succeeding edges both point into it, i.e., the path contains $V_{k-1}\rightarrow V_{k}\leftarrow V_{k+1}$. Given disjoint sets of vertices $A,B$ and $C$, the sets $A$ and $B$ are said to be d-separated given $C$, denoted $A\perp\!\!\!\perp_{\text{d-sep}}B\mid C$, if and only there is no path from a vertex in $A$ to one in $B$ along which (i) every collider on the path is either in $C$ or has a descendant in $C$ and (ii) every non-collider on the path is not in $C$. Paths satisfying conditions (i) and (ii) are said to be open. This definition leads to the following global Markov property of m-DAGs—given disjoint sets $A,B,C$ we have, $A\perp\!\!\!\perp_{\text{d-sep}}B\mid C\implies A\perp\!\!\!\perp B\mid C\text{ in }p(V)$. That is, d-separation in $\mathcal{G}$ implies conditional independences in the probability distribution on the random variables displayed in the m-DAG. This gives us an intuitive way to reason about missingness in FL systems, as we now demonstrate.
@@
However, there is another complication arising from user opt-out that likely results in gradients that are missing not at random (MNAR). The data are MNAR if missingness is not independent of the missing variable given observed covariates alone, i.e., $R\not\perp\!\!\!\perp G(\theta)^{miss}\mid D$ in our FL setup. This occurs when user opt-out is influenced by the data $X,Y$ itself—e.g., a user may not want to share interactions with the model $h_{\theta}$ involving sensitive data $X$ or if they are dissatisfied with model predictions of their outcomes YY—encoded by the edges $X\rightarrow R\leftarrow Y$. This leads to open paths—e.g., $R\leftarrow Y{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\implies}G(\theta)^{miss}$—that imply $R\not\perp\!\!\!\perp G(\theta)^{miss}\mid D$.
@@ 4. Reweighted Device Selection
4:  Prompt all users $u\in\mathcal{U}$ for participation, record $R$
5:  Prompt all users $u\in\mathcal{U}$ for satisfaction, record $S^{miss}$
@@
7:  Define ${\mathcal{U}}_{R}$ as users $u\in\mathcal{U}$ such that $R=1$
8:  for $i$ from 11 to maxiterationsmax\ iterations do
@@
16:    Broadcast $\theta^{(t+1)}$ to all users $u\in\mathcal{U}$
@@ 5. Preliminary Results & Discussion
Now that we have a working prototype implementation, our next step is to run diverse experiments using our Flower implementation of FLOSS. The Flower framework allows us to extensively evaluate the performance of FLOSS using larger datasets in more realistic settings. We will also quantify the computational overhead of Algorithm 1 to show how feasible FLOSS is for real world deployments. Finally, we will experimentally show how FLOSS scales to evaluate the overall generalizability and accuracy of our approach.
6. Conclusion
@@
Acknowledgements.
RB acknowledges support from the NSF CRII grant 2348287. The content of the information does not necessarily reflect the position or the policy of the Government, and no official endorsement should be inferred.